Suppose there is XYZ Company and there are different departments like production, marketing, finance, sales etc. Manager of the company want to know about the detail of the employees who are highly paid in each of the department. Write a program using the concept of classes to implement the same
1
Expert's answer
2021-10-08T23:50:26-0400
class XYZ:
    def __init__(self, department, salary):
        self.department = department
        self.salary = salary
    def __str__(self) -> str:
        return f"Department: {self.department} Salary: {self.salary}\n"


e1 = str(XYZ("production", 120))
e2 = str(XYZ("marketing", 300))
e3 = str(XYZ("finance", 250))
e4 = str(XYZ("sales", 100))
print(e1, e2, e3, e4)

Example:
Department: production Salary: 120
 Department: marketing Salary: 300
 Department: finance Salary: 250
 Department: sales Salary: 100